Solve the following linear equations. 
m – (m - 1)/2 = 1 – (m - 2)/3 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

m–(m-1)/2=1–(m-2)/3
⇒m–m/2–1/2=1–(m/3–2/3)
⇒m–m/2+1⁄2=1–m/3+2/3
⇒m–m/2 + m/3 = 1 + 2/3–1⁄2
⇒m/2+m/3=1⁄2+2/3
⇒(3m+2m)/6=(3+4)/6
⇒5m/6=7/6
⇒m=7/6×6/5
⇒m=7/5

Thanks